7. Bromomethane melts at-93.6 °C. If its 4usHP = 5.98 kJ/mol, determine its Δfuss®.
media%2F064%2F064f5911-b9a1-4aff-a88e-3e
media%2F27c%2F27cac9b6-809f-4823-8a5b-75
Answer #1

\DeltaS fus   =   \DeltaH fus/T

\DeltaH fus     = 5.98Kj/mole    = 5980J/mole

T               = -93.6+ 273   = 179.4K

\DeltaS fus   =   \DeltaH fus/T

                = 5980/179.4    = 33.34J/mole-K >>>>answer